OpenStudy (anonymous):

Water freezes at 0 degrees Celsius and 32 degrees Fahrenheit. Water boils at 100 degrees Celsius and 212 degrees Fahrenheit. A. How would I write degrees Celsius as a linear function of degrees Fahrenheit? B. What is the slope of the linear equation and what does it mean? C. What is the y-intercept and what does it mean?

OpenStudy (anonymous):

with the given numbers how do i create a linear function?

OpenStudy (shamil98):

A. (32,0) and (212,100). M = 100 - 0 /212 - 32 = 100/180 = 5/9 C - 0 =5/9 (F - 32) or C = 5/9(F-32)

OpenStudy (shamil98):

B. The slope is 5/9; it means that that for each one-degree increase in degrees Fahrenheit, degrees Celsius increase by 5/9.

OpenStudy (shamil98):

C. The y intercept is -160/9, and it means that when F is 0 Celsius is at -160/9.

OpenStudy (shamil98):

There ya go.

OpenStudy (anonymous):

Thanks @shamil98 youre the best!!

OpenStudy (shamil98):

I literally did these problems today xD , haha thanks.

OpenStudy (anonymous):

Hahah thats awesome! I just got them today :p

OpenStudy (anonymous):

& no thank you!!!